a 30.0 mL sample of 0.250 M HClO is titrated with 0.125 M KOH. Calculate the pH of the solution after the addition of the following volumes of base. Ka= 3.5 x 10^-8 for HClO.

a) 0.0 mL base added b) 25.0 mL c) 50.0 mL d) 55.0 mL e) 60.0 mL f) 65.0 mL g) 75.0 mL
Answer #1

て-46 (a) o-o ol ef baBe addled 回Noefmmolesebbage = ½(1.46-16(0950)) =て-5 = 4-03 (b) No. etrnmolesekalid=30 xo450 8125 밞 1049